Will you help me find the solution to this equation: 6x-3(6-5x)+3x=10-4(2-x)
I tried this:
6x-18+15x+3x=10-8+4x
21x-15=2+4x
25x-15=2
25x=17
x=17/25
I know the answer is supposed to be 1, I just don't know how to get it.

Answer by stanbon(75887)


6x-18+15x+3x=10-8+4x
24x-18=2+4x
20x=20
x=1
Cheers,
Stan H.